A box contains 18 light bulbs. Of these, four are defective. Five bulbs are chosen at random. What is the probability that four are defectives?


1
Expert's answer
2022-01-26T17:45:44-0500

We have the total of 18bulbs

14 are good and 4 are defectives.

Total number of outcome is "\binom{18}{5}\\"

Selecting 1 good bulb from 14 is "\binom{14}{1}"

Selecting 4 defectives bulb from 4 is "\binom{4}{4}\\"


Then the solution is

"\frac{\binom{14}{1}\times\binom{4}{4}}{\binom{18}{5}}=\frac{14 \times 1}{8568}=0.0016"